# Synthetic Augmentation in Imbalanced Learning: When It Helps, When It Hurts, and How Much to Add

### v1

@@ 7 Simulation Studies
In this section, we use controlled simulations to validate the theory developed above and to evaluate the practical benefits of Validation-Tuned Synthetic Size (VTSS) (Section 6). Our experiments are designed to demonstrate that the key qualitative predictions of our analysis hold in practice: synthetic augmentation helps under local asymmetry, but not under local symmetry. We also show that VTSS more reliably converts augmentation into improved downstream performance than the naive balancing rule $\tilde{n}=n_{0}-n_{1}$.
@@ 7.1 Under Local Asymmetry
We first investigate simulations under local asymmetry, considering both cases with and without directional alignment. We further illustrate the theory under both realistic and inconsistent generators.
Classification Without Directional Alignment.
Figure 2: Balanced test performance under SMOTE with naive balancing versus VTSS.
We simulate an imbalanced binary classification problem in $d=5$ where the majority class is i.i.d. $\bm{x}\mid y=0\sim\mathcal{N}(0,I_{d})$ and the minority class is i.i.d. from a two-component Gaussian mixture: with probability $1/2$, $\bm{x}\mid y=1\sim\mathcal{N}(\delta\bm{e}_{1}+\tfrac{\xi}{2}\bm{e}_{2},I_{d})$ and with probability $1/2$, $\bm{x}\mid y=1\sim\mathcal{N}(\delta\bm{e}_{1}-\tfrac{\xi}{2}\bm{e}_{2},I_{d})$, with $\delta=2.0$ and $\xi=2.5$. Here $\bm{e}_{i}$ denotes the $d$-dimensional standard basis vector. For each scale $m\in\{1,2,4,8,16\}$, we set $n_{1}=50m$ and $n_{0}=20n_{1}$, generate an imbalanced training set, and compare two SMOTE-based training strategies: (A) naive balancing, which oversamples the minority to exactly $n_{0}$ points via SMOTE (with $k=5$ nearest neighbors), and (B) VTSS, which selects $\tilde{n}$ by minimizing balanced validation log-loss over a candidate grid (Section 6). In each case, we fit logistic regression on the SMOTE-augmented training data and evaluate performance on a large balanced test set generated from the true class-conditional distributions, with equal class sizes. Performance is reported using balanced log-loss and balanced test 011 error, averaged over 100 independent repetitions for each $m$.
As shown in Figure 2, both the balanced test log-loss and the balanced test error decrease as $n_{1}$ increases, reflecting the benefit of having more real minority observations. More importantly, VTSS consistently attains lower loss and error than the naive balancing rule and typically exhibits smaller variance, indicating that validation-based tuning yields a more stable procedure by avoiding over- or under-synthesis. The performance gap also narrows as $n_{1}$ grows: when the minority sample becomes sufficiently large, the tuned synthetic size tends to approach the naive target $n_{0}-n_{1}$, so the incremental benefit diminishes. This setting does not exhibit strong directional alignment between the generator mismatch and the intrinsic imbalance effect, so our theory suggests limited gains from fine-tuning beyond near-balancing. In these simulations, VTSS does not underperform the naive choice $\tilde{n}=n_{0}-n_{1}$. In practice, since one rarely knows a priori whether the minority sample size is already large enough for the gap to be negligible, VTSS provides a robust default.
Classification with Directional Alignment.
Figure 3: Balanced excess risk and parameter estimation error in a 2D Gaussian model with aligned synthetic bias.
@@
To illustrate the directional-alignment phenomenon discussed in Section 4.3, we consider a 2D Gaussian setting as in Example 2, with imbalance ratio $n_{0}/n_{1}=20$. For $n_{1}$ varying from 6 to 3200, we set $n_{0}=20n_{1}$, draw majority samples $\bm{x}\mid y=0\sim\mathcal{N}(0,I_{2})$ and minority samples $\bm{x}\mid y=1\sim\mathcal{N}(\bm{\mu}_{1},I_{2})$ with $\bm{\mu}_{1}=(\mu,0)$ and $\mu=1$. We augment the training data with $\tilde{n}$ synthetic points labeled as minority, generated from a biased but directionally aligned generator $\bm{x}_{\mathrm{syn}}\sim\mathcal{N}(\bm{\mu}_{s},I_{2})$ with $\bm{\mu}_{s}=(a,0)$ and $a=0.5$. For each repetition (100 Monte Carlo runs per $n_{1}$), we also generate a fresh balanced test set of size 5000 per class from the true distributions. We compare two fixed synthetic-size rules: the naive balancing choice $\tilde{n}=n_{0}-n_{1}$ and the bias-canceling choice $\tilde{n}=4(n_{0}-n_{1})$ (from the analysis of Example 2). On each augmented training set, we fit the linear predictor $\hat{\bm{\theta}}\in\mathbb{R}^{2}$ by least squares and evaluate (i) parameter error $\|\hat{\bm{\theta}}-\bm{\theta}^{*}\|_{2}$ and (ii) balanced excess risk $\mathcal{R}(\hat{\bm{\theta}})-\mathcal{R}(\bm{\theta}^{*})$, where $\mathcal{R}(\bm{\theta})=\|\bm{\theta}\|_{2}^{2}+\tfrac{1}{2}(\bm{\theta}^{\top}\bm{\mu}_{1}-1)^{2}$ and $\bm{\theta}^{*}=(\mu/(\mu^{2}+2),0)$. Finally, we plot the mean and standard deviation of the excess risk and parameter error versus $n_{1}$ on log scales for the two choices of $\tilde{n}$.
@@
The simulation in Figure 3 confirms the theoretical analysis in Example 2. When $\tilde{n}=4(n_{0}-n_{1})$, both the balanced excess risk and the parameter error $\|\hat{\bm{\theta}}-\bm{\theta}^{*}\|_{2}$ decay to zero as the sample size grows, whereas under naive balancing they converge to a strictly positive level. This demonstrates that count balancing alone can be inconsistent in the presence of biased synthetic data.
@@
Figure 4: Excess risk and parameter error versus minority sample size under a 2D Gaussian mean-shift model (squared loss).
@@
For directional alignment with a realistic and consistent synthetic generator (Section 4.2), we simulate the same 2D Gaussian setup as above under imbalance $n_{0}/n_{1}=20$ and squared loss. For $n_{1}\in[100,5000]$, we run 100 Monte Carlo repetitions, generating a fresh imbalanced training set and a balanced validation set of 1000 points per class. Synthetic minority samples are produced by a realistic, consistent, and directionally aligned generator: for a requested synthetic size $k$, draw $\bm{x}_{\mathrm{syn}}\sim\mathcal{N}(\bm{\mu}_{\mathrm{syn}},I_{2})$ with
$\bm{\mu}_{\mathrm{syn}}=\left(1-(\log n_{1})^{-1/2}\right)\bm{\mu}_{1},$
so that the synthetic mean approaches $\bm{\mu}_{1}$ as $n_{1}$ increases while remaining aligned with $\bm{\mu}_{1}$. We compare VTSS with the naive balancing rule $\tilde{n}=n_{0}-n_{1}$, fit a least-squares predictor on the augmented training data, and evaluate both methods using population excess risk $\mathcal{R}(\hat{\bm{\theta}})-\mathcal{R}(\bm{\theta}^{*})$ and parameter error $\|\hat{\bm{\theta}}-\bm{\theta}^{*}\|_{2}$, reporting mean $\pm$ standard deviation over repetitions as functions of $n_{1}$.
@@
The results are shown in Figure 4. Across all values of $n_{1}$, VTSS achieves substantially smaller errors and a faster decay rate, supporting our theory that, under a realistic but consistent generator, directional information can be exploited to select $\tilde{n}$ that outperforms naive balancing. In contrast to Figure 2, where the gap between the two methods diminishes as $n_{1}$ grows, here the gap remains substantial, highlighting the effect of directional alignment.
@@ 7.2 Under Local Symmetry
Figure 5: Balanced excess risk $\mathcal{R}(\hat{\bm{\theta}})-\mathcal{R}(\bm{\theta}^{*})$ (log scale) versus the synthetic-size multiplier $\gamma$.
To further illustrate local symmetry (Section 5), we conduct a simulation study based on the mean-shift model in Example 3. We generate imbalanced binary data in dimension $d=20$ from
$\bm{x}\mid y=1\sim\bm{\mu}+\bm{\xi},\qquad\bm{x}\mid y=0\sim-\bm{\mu}+\bm{\xi},$
[…]
The results in Figure 5 align with the theoretical analysis. In the mean-shift setting, the majority–minority difference cancels at the balanced optimum ($\nabla\phi(\bm{\theta}^{*})=0$), so adding synthetic data from a realistic generator cannot reduce this part of the error. Instead, any mismatch between the synthetic and true minority distributions introduces bias and can only worsen balanced excess risk. Empirically, realistic generators, including the Semi-Oracle generator that is already close to the truth, fail to improve performance and typically degrade it as more synthetic samples are added. In contrast, only the Oracle generator improves monotonically with increasing $\gamma$, reflecting that it avoids generator-induced mismatch and helps by adding correctly distributed minority samples.
To further demonstrate a regime where class imbalance is not the bottleneck, we consider the sigmoid Bernoulli logistic model in Example 4 with dimension $d=20$. We let $v=\bm{e}_{1}$ and $\bm{\theta}_{\mathrm{true}}=c\bm{v}$ with $c=1$, and generate covariates as $\bm{x}=T\bm{v}+W$, where $T\in\{-a,b\}$ with $a=5$, $b=1$, and $\Pr(T=-a)=\alpha$ chosen to ensure bias cancellation. The noise $W$ is orthogonal to $\bm{v}$ (its first coordinate is identically zero) and has a symmetric, non-Gaussian mixture distribution in the remaining coordinates: $W_{2:d}=\bm{z}+s\bm{\mu}$ with $\bm{z}\sim\mathcal{N}(0,I)$, $s\in\{\pm 1\}$ equally likely, and $\bm{\mu}$ having its first three entries equal to $2.5$ (others zero), making Gaussian-fit imperfect. Labels are drawn as $y\in\{0,1\}$ with $\Pr(y=1\mid\bm{x})=\sigma(\bm{\theta}_{\mathrm{true}}^{\top}\bm{x})$. For each of 100 repetitions, we sample $n_{\mathrm{train}}=5000$ i.i.d. training pairs $(\bm{x},y)$ (with random realized class counts), define the minority class as the label with fewer observations, and augment only that minority class to size $n_{1}+\tilde{n}$, where $\tilde{n}=\mathrm{round}(\gamma(n_{0}-n_{1}))$ and $\gamma$ ranges over $[0,2]$. We compare three generators for minority augmentation: SMOTE (nearest-neighbor interpolation with $k=5$), Gaussian-fit (sampling from a fitted Gaussian using the minority empirical mean and covariance with a small ridge), and an Oracle generator that draws fresh $(\bm{x},y)$ from the true model and keeps only minority-labeled points. For each augmented dataset, we fit logistic regression and report (i) balanced excess logistic risk relative to $\bm{\theta}^{*}$ and (ii) parameter error $\|\hat{\bm{\theta}}-\bm{\theta}^{*}\|_{2}$, where $\bm{\theta}^{*}$ is obtained by fitting the same model on a large balanced sample of 20,000 per class.
@@ 7.3 Further Evaluation of VTSS
In this subsection, we further illustrate the advantage of VTSS by plotting the test loss as a function of the synthetic size $\tilde{n}$ and comparing it with the loss achieved by VTSS. We consider two configurations.
Linear classification (SMOTE + logistic regression).
Nonlinear classification (perturbed sampling + kernel logistic).
We use a non-Gaussian data-generating process in $d=5$. For the majority class $y=0$, we first draw a mixture indicator $z\sim\mathrm{Bernoulli}(0.30)$, then sample the first two coordinates from a correlated Gaussian $(x_{1},x_{2})\sim\mathcal{N}\left((1.5\cdot z,0),\begin{pmatrix}1&0.5\\ 0.5&1\end{pmatrix}\right)$, and sample the remaining coordinates i.i.d. as $x_{3:5}\sim\mathcal{N}(0,I_{3})$. For the minority class $y=1$, we generate a noisy ring in the $(x_{1},x_{2})$-plane by drawing an angle $\omega\sim\mathrm{Unif}[0,2\pi]$ and radius $r=2.2+\varepsilon_{r}$ with $\varepsilon_{r}\sim\mathcal{N}(0,0.12^{2})$, then setting
$x_{1}=\delta+r\cos\omega+\varepsilon_{1},\qquad x_{2}=r\sin\omega+\varepsilon_{2},$
with $\varepsilon_{1},\varepsilon_{2}\sim\mathcal{N}(0,0.08^{2})$ and $\delta=4.8$. The remaining coordinates are small Gaussian noise $x_{3:5}\sim\mathcal{N}(0,0.08^{2}I_{3})$. This simulation is a controlled nonlinear-boundary benchmark, designed to test whether VTSS can tune synthetic size when the generator introduces distributional blur. We embed the 2D structure in $d=5$ by adding nuisance coordinates, reflecting the common setting where only a subset of features drives class separation.
Synthetic data are generated by perturbed resampling: bootstrap minority points and add Gaussian jitter $\mathcal{N}(0,I)$. We train an RBF-kernel probabilistic classifier, sweep $\tilde{n}=\mathrm{round}(\gamma(n_{0}-n_{1}))$ over $\gamma\in[0.6,1.4]$, and let VTSS pick $\tilde{n}$ via validation minimization in each repetition.
@@
In both configurations of Figure 8, the empirical curves confirm the theoretical insight that, although the naive balancing choice $\tilde{n}=n_{0}-n_{1}$ often lies near the optimum, it is generally not the loss-minimizing synthetic size. Moreover, VTSS, which selects $\tilde{n}$ by minimizing balanced validation loss on each dataset, achieves substantially lower average loss than any single fixed choice of $\tilde{n}$ over 100 repetitions. This highlights the benefit of data-dependent tuning around a theory-guided target.

### v2

@@ 7 Simulation Studies
In this section, we use controlled simulations to validate the theory developed above and to evaluate the practical benefits of Validation-Tuned Synthetic Size (VTSS) (Section 6). Due to space constraints, we defer the detailed simulation study settings to Section 15 of the Supplement.
@@ 7.1 Under Local Asymmetry
In the local asymmetry case, we study the classification with directional alignment. To illustrate the directional-alignment phenomenon for biased synthetic generator discussed in Section 4.3, we consider a 2D Gaussian setting as in Example 2, with a linear predictor. We compare two fixed synthetic-size rules: the naive balancing choice $\tilde{n}=n_{0}-n_{1}$ and the bias-canceling choice $\tilde{n}=4(n_{0}-n_{1})$ (from the analysis of Example 2). We evaluate parameter error $\|\hat{\bm{\theta}}-\bm{\theta}^{*}\|_{2}$ and balanced excess risk.
Figure 2: Balanced excess risk and parameter estimation error in a 2D Gaussian model with aligned synthetic bias based on $100$ simulations.
@@
The simulation in Figure 2 confirms the theoretical analysis in Example 2. When $\tilde{n}=4(n_{0}-n_{1})$, both the balanced excess risk and the parameter error $\|\hat{\bm{\theta}}-\bm{\theta}^{*}\|_{2}$ decay to zero as the sample size grows, whereas under naive balancing they converge to a strictly positive level. This demonstrates that naive balance can be inconsistent in the presence of biased synthetic data.
@@
Figure 3: Excess risk and parameter error versus minority sample size under a 2D Gaussian mean-shift model (squared loss) based $100$ simulations.
@@
For directional alignment with a realistic and consistent synthetic generator (Section 4.2), we simulate the same 2D Gaussian setup with squared loss. Synthetic minority samples are produced by a realistic, consistent, and directionally aligned generator: for a requested synthetic size $k$, draw $\bm{x}_{\mathrm{syn}}\sim\mathcal{N}(\bm{\mu}_{\mathrm{syn}},I_{2})$ with $\bm{\mu}_{\mathrm{syn}}=\left(1-(\log n_{1})^{-1/2}\right)\bm{\mu}_{1},$ so that the synthetic mean approaches $\bm{\mu}_{1}$ as $n_{1}$ increases while remaining aligned with $\bm{\mu}_{1}$. We compare VTSS with the naive balancing rule $\tilde{n}=n_{0}-n_{1}$, and evaluate population balanced excess risk and parameter error.
@@
The results are shown in Figure 3. Across all values of $n_{1}$, VTSS achieves substantially smaller errors and a faster decay rate, supporting our theory that, under a realistic but consistent generator, directional information can be exploited to select $\tilde{n}$ that outperforms naive balancing.
@@ 7.2 Under Local Symmetry
Figure 4: Balanced excess risk $\mathcal{R}(\hat{\bm{\theta}})-\mathcal{R}(\bm{\theta}^{*})$ (log scale) versus the synthetic multiplier $\gamma$.
To further illustrate local symmetry (Section 5), we conduct a simulation study based on the mean-shift model in Example 3 with a linear predictor. We consider three choices of $\bm{\xi}$: Uniform cube, Rademacher, and Uniform sphere. We start from an imbalanced training sample with $n_{1}=100$ and $n_{0}=2000$, and augment the minority class with $\tilde{n}=\mathrm{round}(\gamma(n_{0}-n_{1}))$ synthetic samples, where $\gamma$ ranges over 20 values in $[0,4]$. We compare four synthetic generators: (i) Oracle; (ii) SMOTE; (iii) Gaussian-fit; and (iv) Semi-Oracle, which samples $\bm{\xi}$ from the true noise distribution but centers at the empirical minority mean $\hat{\bm{\mu}}_{1}$ (i.e., $\hat{\bm{\mu}}_{1}+\bm{\xi}$).
The results in Figure 4 align with the theoretical analysis. In the mean-shift setting, the majority–minority difference cancels at the balanced optimum ($\nabla\phi(\bm{\theta}^{*})=0$), so adding synthetic data from a realistic generator cannot reduce this part of the error. Instead, any mismatch between the synthetic and true minority distributions introduces bias and can only worsen balanced excess risk. Empirically, realistic generators, including the Semi-Oracle generator that is already close to the truth, fail to improve performance and typically degrade it as more synthetic samples are added. In contrast, only the Oracle generator improves monotonically with increasing $\gamma$, reflecting that it avoids generator-induced mismatch and helps by adding correctly distributed minority samples. For the VTSS, we plot the empirical distribution of the selected $\gamma^{*}$ in Figure 9 of the Supplementary Materials. VTSS selects a very small $\gamma^{*}$ in the vast majority of runs and frequently chooses $\gamma^{*}=0$ exactly, further demonstrating the effectiveness of the proposed VTSS framework.
We also have analogous local symmetry results showing that synthetic data provides no benefit for the sigmoid Bernoulli logistic model as in Example 4 these results are presented in Section 15.2.2 of the supplementary materials.
@@ 7.3 Further Evaluation of VTSS
In this subsection, we further illustrate the advantage of VTSS by plotting the test loss as a function of the synthetic size $\tilde{n}$ and comparing it with the loss achieved by VTSS. We consider two configurations. (i) Linear classification (SMOTE + logistic regression) and (ii) Nonlinear classification (perturbed sampling + kernel logistic).
@@
In both configurations of Figure 5, the empirical curves confirm the theoretical insight that, although the naive balancing choice $\tilde{n}=n_{0}-n_{1}$ often lies near the optimum, it is generally not the loss-minimizing synthetic size. Moreover, VTSS, which selects $\tilde{n}$ by minimizing balanced validation loss on each dataset, can always capture the lowest average loss. This highlights the benefit of data-dependent tuning around a theory-guided target.
8 Real Data Applications
In this section, we present real data applications to further illustrate our theoretical findings and evaluate the proposed VTSS method.
Data source, cohort construction, features.
We study prediction of sepsis-related outcomes using the MIMIC-III critical care database. Clinically, sepsis is a life-threatening condition arising from a dysregulated response to infection that can rapidly progress to organ dysfunction, making timely identification essential for improving outcomes (Singer et al. 2016). We conduct experiments on MIMIC-III (Medical Information Mart for Intensive Care), a large, de-identified, single-center ICU dataset covering admissions at Beth Israel Deaconess Medical Center from 2001–2012, with rich clinical information including demographics, bedside vital signs, laboratory results, medications, clinical notes, and mortality outcomes (Johnson et al. 2016).
Using MIMIC-III, we construct an adult admission cohort (age >15>15 years at admission). We consider three binary prediction tasks, with in-hospital mortality, sepsis, and septic shock as responses. These represent clinically consequential outcomes capturing (i) overall patient prognosis, (ii) infection-related critical illness, and (iii) progression to the most severe end of the sepsis spectrum. Sepsis status is identified using ICD-9 diagnosis codes, labeling admissions with severe sepsis or septic shock as sepsis cases.
To capture early physiologic trajectories, we extract 10 clinical variables measured during the first 120 hours after admission and use binned averages to stabilize irregularly sampled measurements:
  • Bedside vital signs from CHARTEVENTS, averaged into 6-hour bins (20 bins): heart rate, respiratory rate, temperature, and systolic blood pressure.
  • Laboratory measurements from LABEVENTS, averaged into 24-hour bins (5 bins): white blood cell count, platelets, creatinine, INR, lactate, and total bilirubin.
For each admission, we convert the multivariate time series into a mean feature vector over the observation window. Missing features are imputed using the corresponding cohort-level feature means. After preprocessing, all three datasets are substantially imbalanced; the class counts and imbalance ratios are summarized in Table 1.
Positive Samples ($n_{1}$) Negative Samples ($n_{0}$) Imbalance Ratio ($n_{0}/n_{1}$)
Mortality 5,7895{,}789 45,06845{,}068 $7.785$
Sepsis 4,0844{,}084 46,77346{,}773 $11.45$
Septic Shock 2,5852{,}585 48,27248{,}272 $18.67$
Table 1: Class imbalance for the mortality, sepsis, and septic shock prediction tasks.
We train (i) logistic regression and (ii) support vector machine (SVM) classifiers on these features to predict each outcome.
Evaluation metric: balanced excess risk.
Performance is primarily assessed using the balanced excess risk. Because the population minimizer $\theta^{*}=\arg\min_{\theta}\mathcal{R}(\theta)$ depends on unknown population expectations, we approximate the subtractive term using an oracle plug-in estimate on each test split. Specifically, we compute $\theta^{*}_{\text{test}}=\arg\min_{\theta}\hat{\mathcal{R}}_{\text{test}}(\theta)$ (the parameter that minimizes the balanced loss on the test set) and report the balanced excess risk as $\hat{\mathcal{R}}_{\text{test}}(\hat{\theta})-\hat{\mathcal{R}}_{\text{test}}(\theta^{*}_{\text{test}})$. To account for sampling variability, we repeat the train-test split 100 times with different random seeds and report the average balanced excess risk across repetitions.
Synthetic data generation and tuning protocol.
We compare four synthetic data generators: SMOTE, ADASYN, Borderline-SMOTE, and Minority Jitter (randomly sampling a minority instance and adding feature-wise Gaussian perturbations). Let $\gamma$ denote the synthetic size multiplier. For each task–classifier–generator combination, we tune $\gamma$ via 5-fold cross-validation and repeat the cross-validation procedure 10 times. We sweep $\gamma$ over a range centered around 1 and report the mean balanced excess risk with a 95% confidence interval at each $\gamma$. For reference, we also plot a horizontal dashed line (with 95% confidence interval) corresponding to the average balanced excess risk achieved by VTSS.
Raw-Mortality
0.3825[0.3809,0.3842]0.3825~[0.3809,0.3842]
Raw-Sepsis
0.5721[0.5700,0.5742]0.5721~[0.5700,0.5742]
Raw-Septic shock
0.7786[0.7761,0.7810]0.7786~[0.7761,0.7810]
Figure 6: Excess risk in predicting hospital mortality, sepsis and septic shock across four synthetic generators using logistic regression. Bottom row is the excess risk obtained by training with raw imbalanced data without synthetic data augmentation.
Results for logistic regression.
Figure 6 highlights that the optimal synthetic sample size is highly sensitive to both the clinical endpoint and the chosen generator. Training on the raw imbalanced data without any synthetic augmentation (corresponding to the $\gamma=0$ case) can yield substantially worse performance, as suggested the bottom row of Figure 6, underscoring the need to address class imbalance in these tasks. Across mortality, sepsis, and septic shock, the balanced excess risk curves as a function of the synthetic multiplier $\gamma$ exhibit minima at markedly different locations: depending on the task and generator, the best-performing $\gamma$ can be less than 1 (conservative oversampling), close to 1, or greater than 1 (aggressive oversampling). This variability indicates that naively fixing $\gamma=1$ may incur a nontrivial excess-risk penalty in some settings. In contrast, VTSS provides a stable and reliable choice for selecting synthetic size. Its performance matches (within uncertainty) the best achievable risk across the candidate $\gamma$ values considered, effectively recovering the balanced excess risk at optimal synthetic size. Therefore, VTSS can be used as a robust default when choosing the synthetic multiplier across different prediction tasks and generators.
Raw-Mortality
0.1463[0.1428,0.1497]0.1463~[0.1428,0.1497]
Raw-Sepsis
0.2138[0.2091,0.2185]0.2138~[0.2091,0.2185]
Raw-Septic shock
0.2478[0.2427,0.2529]0.2478~[0.2427,0.2529]
Figure 7: Excess risk in predicting hospital mortality, sepsis and septic shock across four synthetic generators using SVM. Bottom row is the excess risk obtained by training with raw imbalanced data without synthetic data augmentation.
Results for SVM.
We next repeat the same experimental pipeline under the identical framework, replacing logistic regression with an SVM classifier. The results are summarized in Figure 7. The qualitative conclusions remain consistent: performance is sensitive to $\gamma$, the optimal $\gamma$ depends on both the endpoint and generator, and training on the raw imbalanced data is frequently suboptimal according to the bottom row of Figure 7. Notably, for in-hospital mortality the risk-minimizing multipliers are often concentrated around $\gamma\in[0.4,0.6]$, suggesting that under SVM this task favors substantially less oversampling than the naive balancing choice $\gamma=1$. VTSS captures this behavior and consistently recovers near-minimal balanced excess risk, supporting its use as a robust default across classifiers, tasks, and generators.
Alternative validation objective: balanced accuracy.
To further assess the flexibility of VTSS, we also consider an alternative validation objective motivated by Remark 2. Instead of selecting $\gamma$ by minimizing validation balanced log-loss, we instantiate VTSS with a decision-oriented criterion and choose $\gamma$ to maximize validation balanced accuracy, keeping all other components of the pipeline unchanged. Figure 8 reports balanced accuracy as a function of $\gamma$ for each generator, along with the VTSS-selected operating point and associated uncertainty.
Figure 8 shows that VTSS consistently selects $\gamma$ values achieving near-optimal balanced accuracy across generators, while the naive balancing choice can be noticeably suboptimal. We also observe that VTSS does not always coincide exactly with the global maximizer of the balanced accuracy curve. A plausible explanation is that balanced accuracy is a thresholded, non-smooth metric that can exhibit flat or step-like validation profiles across nearby $\gamma$ values, making the exact maximizer difficult to localize under finite-sample variability. Nevertheless, VTSS remains consistently close to the best achievable balanced accuracy while improving upon naive balancing, supporting its use as a robust default for synthetic-size selection under alternative evaluation metrics.
Figure 8: Balanced accuracy in predicting hospital mortality across four synthetic generators.